# Uncertainty-Aware Diffusion Guided Refinement of 3D Scenes

### v1

@@ 4 Experiments
We evaluate UAR-Scenes on the task of Novel View Synthesis (NVS) of scenes. Keeping in mind that multi-view consistency is a major focus to assess the quality of refined scenes, we present results on both narrow and wide baselines in Table 1. We further evaluate the interpolation and extrapolation performance following [1] and [30] after refinement in Table 2. To highlight the transferability of our method across datasets, we also show NVS results on KITTI-v2 in Table 3. This shows our refinement pipeline can be seamlessly integrated with any existing reconstruction method across a wide variety of tasks. Finally, we present ablation studies in Section 5.3 showing the importance of each component in our framework and how it contributes towards overall performance.
Figure 4: Qualitative Results. (a) Qualitative comparisons on the RealEstate-10K dataset shows that our method produces more realistic results which are more plausible and faithful to the original image (In 2nd row, 3rd column, UAR-Scenes is able to render leg of table which is blurry in Flash3D’s prediction). (b) Qualitative comparison on the KITTI-v2 dataset which shows that our method can deliver sharp results especially in edges where there may be ambiguity (Back edge of car is distorted in Flash3D’s prediction in 2nd column). Notice that despite the significant camera motion between the original input view and the target novel views, UAR-Scenes can render realistic and plausible renderings as highlighted above.
@@ 4.1 Experimental Setup
Evaluation Metrics. We adopt PSNR, SSIM and Perceptual Similarity (LPIPS) as our photo-metric evaluation standard as in [11, 12, 1]. Further, we report Frechet-Inception Distance (FID) [58] as well since we generate plausible explanations for unobserved regions which do not have a corresponding paired ground truth.
@@
Optimization Details. We optimize each scene for 1000 steps in which the learning rate of position information decays from $1\times 10^{-3}$ to $2\times 10^{-5}$. We use a single NVIDIA L40 48 GB GPU for running all our experiments. Finally, our rendering resolution for all images is $256\times 384$ following the protocol in [1]. We keep a batch size of 2 for all our experiments. More implementation details are listed in the supplementary material.
4.2 Novel View Synthesis.
@@
4.3 Generative Results.
We report interpolation and extrapolation results on the RealEstate-10K dataset using Flash3D as the baseline method. We are able to beat all the methods except MVSplat which uses stereo information to interpolate. In Extrapolation, over a wide baseline, our FID is significantly lower then all methods including LatentSplat [30] which uses a GAN denoiser but fails to handle complex real-world scenes.
@@ 5 Conclusion
We introduced UAR-Scenes, a novel 3D scene refinement pipeline that enhances scene Gaussians derived from a single image, thereby improving the quality of Gaussians produced by 3D reconstruction pipelines. Our results demonstrate the versatility of UAR-Scenes across both in-domain and out-domain datasets. We also outperform state-of-the-art feed-forward methods across small, medium, and large baseline settings. Additionally, UAR-Scenes excels in challenging interpolation and extrapolation tasks, yielding superior rendered views and high-quality generations at unseen camera angles, as evidenced by both qualitative and quantitative results. Our ablation studies validate the effectiveness of the components and design choices within the UAR-Scenes pipeline, highlighting the benefits of employing Fourier-style texture alignment with real-world scenes. Thus we highlight UAR-Scenes’s role to advance 3D scene understanding and novel view synthesis.
Figure 6: Uncertainty Map Pipeline. We pass the pseudo views to the MLLM [23] first using the one shot in-context learning setup as shown above. This gives the object tags which is then passed onto the open-vocabulary segmentation model LSeg ($\mathcal{S}$[24]. We then compute the per-pixel entropy obtained from the segmentation maps to generate the corresponding uncertainty maps $\mathcal{U}$.
6 Preliminaries
6.1 3D Gaussian Splatting
Gaussian Primitives ($\gamma_{n}$). A 3D scene can be explicitly represented by a set of anisotropic Gaussian ellipsoids with positions $\mu_{n}\in\mathbb{R}^{3}$ , covariance matrix $\Sigma_{n}\in\mathbb{R}^{3\times 3}$, color $c_{n}\in\mathbb{R}^{3\times(k+1)^{2}}$ for order $k$, typically represented using Spherical Harmonic (SH) coefficients and opacity $\alpha_{n}\in[0,1]$. For each Gaussian point x, it’s 3D position is given as
$G(x)=e^{-\frac{1}{2}\,(x-\mu_{n})^{\top}\,\Sigma_{n}^{-1}\,(x-\mu_{n})},$ (7)
The $\Sigma_{n}$ is decomposed into two learnable components represented by the scaling matrix $S_{n}$ and a rotation matrix $R_{n}$ as follows:
$\Sigma_{n}=R_{n}\,S_{n}\,S_{n}^{\top}\,R_{n}^{\top}.$ (8)
Therefore any scene can be represented as a collection of Gaussian primitives where each primitive can be represented as $\gamma_{n}:=(\mu_{n},R_{n},S_{n},\alpha_{n})$.
Rasterization. The trainable parameters acquired within the primitive $\gamma_{n}$ can be optimized via the application of the ensuing differentiable rendering function:
$I_{0}(p)=\sum_{n=1}^{N}c_{n}\,\tilde{\alpha}_{n}\prod_{m=1}^{n-1}\bigl{(}1-\tilde{\alpha}_{m}\bigr{)},$ (9)
where $I_{0}(p)$ represents the rendered color at pixel p in rendered image $I_{0}$ and $\tilde{\alpha}_{n}$ is calculated from the back-projected 2D Gaussians.
6.2 Latent Video Diffusion Models (LVDMs)
Video Diffusion Model. Latent Video Diffusion Models consist of a pre-trained encoder $\mathcal{E}$, a U-Net denoiser $\epsilon_{\theta}$ and a pre-tained decoder $\mathcal{D}$. The diffusion process occurs in the latent space. Given an image $\mathcal{I}$, it is initially embedded in the latent space via the frozen encoder $\mathcal{E}$ yielding latent $z_{0}^{1:M}=\mathcal{E}\bigl{(}x_{0}^{1:M}\bigr{)}$ by progressively sampling noise from a Gaussian distribution $\epsilon\sim\mathcal{N}(0,I)$ to produce noise $z_{T}$ over T progressive timesteps. This could be given by the equation:
$z_{t}^{1..M}=\sqrt{\overline{\alpha}_{t}}\,z_{0}^{1..M}+\sqrt{1-\overline{\alpha}_{t}}\,\epsilon_{t}^{1..M},$ (10)
where $\alpha_{t}\in(0,1)$, and $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. The denoiser $\epsilon_{\theta}$ is then trained by minimizing the reconstruction loss:
$\mathbb{E}_{x_{0}^{1..M},\,y_{t}^{1..M},\,\epsilon_{t}^{1..M}\sim\mathcal{N}(\textbf{0},I)}\Bigl{\|}\epsilon_{t}^{1..M}-\epsilon_{\theta}\bigl{(}z_{t}^{1..M},t,y\bigr{)}\Bigr{\|}_{2},$ (11)
where $y$ is the input conditioning signal. This trained denoiser can then be used to generate a sequence of M images $I_{1...M}$ given a conditioning image $\mathcal{I}$ at the test time.
6.3 Training Details
Pseudo View Pre-Processing. To prepare the pseudo views, we generate 14 frames using MotionCtrl [22] in both the forward and backward directions and continue to progressively do so until paired pseudo views have been generated for all the frames corresponding to each particular scene in RealEstate-10k. For the out-domain KITTI-v2 dataset, since it follows a stereo format, we generate the pseudo views for the right camera following the standard protocol followed by existing works [16, 1] all of which reconstruct the scene based on views obtained from the left camera and test on novel views from the right camera. We keep the standard test resolution of $375\times 1242$ and crop the outer 5% from all the images following the baseline [1] protocol.
LVDM Details. To denoise the pseudo views, we perform 50 denoising inference steps per image. We follow the same resolution of $256\times 384$ in RealEstate-10K to ensure that the generated images are consistent with the rendered images from the 3D scene. We keep a FPS of 6 for all our experiments and set speed=1. We don’t utilize the motion_bucket_id parameter since it is irrelevant for our use case.
Figure 7: Additional Qualitative Results. (a) UAR-Scenes is seamlessly able to adapt to indoor and outdoor scenes while preserving realistic and plausible quality in areas where Flash3D fails. In some cases as in the 2nd row, our method produces plausible explanations for regions outside of the input image’s view but which may not align with the ground truth image. The FID metric is crucial to assess the effectiveness of our method in such cases.(b) UAR-Scenes similarly generalizes to KITTI as well showing robust performance in previously unseen scenarios.
6.4 MLLM for Object Tagging
In this work, we utilize the BLIP2-Tran-T5-XL [23] model for extracting both partially and fully visible objects. To ensure that the MLLM adheres to the task in hand, we leverage a one-shot in-context learning setup [59] which significantly enhances it’s ability to detect objects. The prompting regimen which we followed for generating the object tags is described in Figure 6.
6.5 Qualitative Results
We present additional qualitative results on the RealEstate-10K [26] and KITTI-v2 [27] datasets shown in Figure 7. As it can be seen, UAR-Scenes shows robust performance across a wide variety of indoor and outdoor scenes. Further, UAR-Scenes is able to produce meaningful explanations outside of the input image’s view owing to it’s strong extrapolation capabilities. Hence, when combined with an existing 3D reconstruction pipeline, we can refine the coarse Gaussians and get more realistic and plausible renderings over a wide variety of real-world scenes.

### v2

@@ 4 Experiments
We evaluate UAR-Scenes on the task of Novel View Synthesis (NVS) of scenes. Keeping in mind that multi-view consistency is a major focus to assess the quality of refined scenes, we present results on both narrow and wide baselines in Table 1. We further evaluate the interpolation and extrapolation performance following [1] and [30] after refinement in Table 2. To highlight the transferability of our method across datasets, we also show NVS results on KITTI-v2 in Table 3. This shows our refinement pipeline can be seamlessly integrated with any existing reconstruction method across a wide variety of tasks. Finally, we present ablation studies in showing the importance of each component in our framework and how it contributes towards overall performance.
Figure 4: Qualitative Results. (a) Qualitative comparisons on the RealEstate-10K dataset shows that our method produces more realistic results which are more plausible and faithful to the original image (In 1st row, Flash3D’s renderings are blurry outside the camera’s seen frustum where UAR-Scenes is able to complete the window. Similarly UAR-Scenes can provide reasonable completions which may not always align with the GT as shown in 2nd row). (b) Qualitative comparison on the KITTI-v2 dataset which shows that our method can deliver sharp results especially in edges where there may be ambiguity (Back edge of car is distorted in Flash3D’s prediction in 2nd column). Notice that despite the significant camera motion between the original input view and the target novel views, UAR-Scenes can render realistic and plausible renderings as highlighted above.
@@ 4.1 Experimental Setup
Evaluation Metrics. We adopt PSNR, SSIM [53] and Perceptual Similarity (LPIPS) [60] as our photo-metric evaluation standard as in [11, 12, 1]. Further, we report Frechet-Inception Distance (FID) [61] as well since we generate plausible explanations for unobserved regions which do not have a corresponding paired ground truth.
@@
Optimization Details. We optimize each scene for 1000 steps in which the learning rate of position information decays from $1\times 10^{-3}$ to $2\times 10^{-5}$. We use a single NVIDIA L40 48 GB GPU for running all our experiments. Finally, our rendering resolution for all images is $256\times 384$ following the protocol in [1]. We keep a batch size of 2 for all our experiments. Further implementation details are listed in the supplementary material.
4.2 Novel View Synthesis
@@
4.3 Generative Results
We report interpolation and extrapolation results on the RealEstate-10K dataset using Flash3D as the baseline method. We are able to beat all the methods except MVSplat which uses stereo information to interpolate. In Extrapolation, over a wide baseline, our FID is significantly lower then all methods including LatentSplat [30] which uses a generative GAN denoiser but still fails to handle complex real-world scenes. Except Flash3D [1], we report FID numbers following existing feed forward methods [30, 62].
@@ 5 Conclusion
We introduced UAR-Scenes, a novel 3D scene refinement pipeline that enhances scene Gaussians derived from a single image, thereby improving the quality of Gaussians produced by 3D reconstruction pipelines. Our results demonstrate the versatility of UAR-Scenes across both in-domain and out-domain datasets. Through Novel View Synthesis experiments, we outperform state-of-the-art feed-forward methods across small, medium, and large baseline settings. Additionally, UAR-Scenes excels in challenging interpolation and extrapolation tasks, yielding superior rendered views and high-quality generations at unseen camera angles, as evidenced by both qualitative and quantitative results. Furthermore, our ablation studies validate the effectiveness of individual components and design choices within the UAR-Scenes pipeline, highlighting the benefits of employing Fourier-style texture alignment with real-world scenes. Overall, our findings highlight UAR-Scenes’s potential to advance 3D scene understanding and novel view synthesis.
Acknowledgments: We gratefully acknowledge the support of USDA NRI grant 2021-67022-33453, UC MRPI grant A21-0101-S003, and NSF grant CMMI-2326309. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.